A plane travels passing coordinate (2,-1,-3) and (4,2,7). At the time of travel, wind

blows in parallel associating to the vector of ."i+ 4j- 2k." Describe the situation

according to the suitable mathematical concepts.

<10 marks>


1
Expert's answer
2021-04-15T07:34:04-0400



Points are

(2,-1,-3) , (4,2,7)



(4-2)i + (2+1)j + (7+3)k


a = "2i + 3j + 10k"

This is vector equation



Wind - vector equation

"b = i+ 4j- 2k"


And

We know


"a.b = |a||b|cos(\theta)"


"cos(\theta) = {a.b \over |a||b|}"


"cos(\theta) = { -6\over \sqrt {113} \sqrt{21}}"


The angle between the plane direction and the wind direction

"\boxed{(\theta) = cos^{-1}({ -6\over \sqrt {113} \sqrt{21}})}"